OpenStudy (anonymous):

PLZZ HELP ILL FAN Which of the following are right triangles? (Choose all that apply.) Question 10 options: a) A triangle with sides of lengths 20 cm, 48 cm, and 50 cm. b) A triangle with sides of lengths 15 mm, 20 mm, and 25 mm. c) A triangle with sides of lengths 21 mm, 72 mm, and 75 mm. d) A triangle with sides of lengths 6 cm, 9 cm, and 12 cm. e) A triangle with sides of lengths 16 in, 32 in, and 34 in. f) A triangle with sides of lengths 9 in, 40 in, and 41 in.

OpenStudy (anonymous):

a) you need to check if 20^2+48^2 = 50^2. if it is, then it's right triangle; You do the same with all other examples

OpenStudy (anonymous):

let's say that "a" is smallest side, "b" medium, and "c" largest, formula is a^2+b^2=c^2; if it's true, then it's right triangle, if not, then it isn't example 1. 3cm, 4cm and 5cm a=3, b=4, c=5 3^2+4^2=5^2 9+16=25 25=25 so it's right triangle but 2. 4cm, 5cm, 6cm 4^2+5^2=6^2 16+25=36 41=36 is not true, so, the triangle is not right hope this helps

OpenStudy (anonymous):

its b, c and f

OpenStudy (anonymous):

yes :)